Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability.

Expert replies
by winniethepooh » Wed Jun 15, 2011 11:08 am
A fair die is rolled once ana a fair coin is flipped once. What is the probability that either the die will land on 3 or that the coin will land on heads?
Join the discussion
Source: — Problem Solving |

by Ashley@VeritasPrep » Wed Jun 15, 2011 12:33 pm
Hey there,

The quickest way to do a problem like this is to add the separate probabilities (of a 3 and of heads) and subtract the overlap once, since we'll have counted that overlap twice in adding. The probability of getting a 3 is 1/6 and the probability of getting heads is 1/2, but that 1/6 includes both head-getting and tail-getting (and the 1/2 includes both 3-getting and non-3-getting). We will get heads 1/2 of the time that we get a 3, so in other words 1/2 * 1/6 of the time. There's your area of overlap.

So your equation becomes 1/6 + 1/2 - (1/6)(1/2). (Note that this method still counts that intersection in which we get a 3 AND heads, but it properly counts it just once instead of twice. If we had simply added 1/6 and 1/2, we would have been counting the intersection twice.)

Does that make sense? If not, I'll clarify with diagrams :)
Ashley Newman-Owens
GMAT Instructor
Veritas Prep

Post helpful? Mosey your cursor on over to that Thank button and click, please! I will bake you an imaginary cake.
Join the discussion

by winniethepooh » Wed Jun 15, 2011 3:27 pm
It does make sense, But I didn't get the perfect picture of the solution!
Join the discussion

by Ian Stewart » Wed Jun 15, 2011 3:55 pm
winniethepooh wrote:A fair die is rolled once ana a fair coin is flipped once. What is the probability that either the die will land on 3 or that the coin will land on heads?
The probability that we don't get a 3 on the die is 5/6, and the probability we don't get Heads is 1/2, so the probability neither event happens is (1/2)(5/6) = 5/12. Thus the probability one or both events happen is 1 - 5/12 = 7/12.
For online GMAT math tutoring, or to buy my higher-level Quant books and problem sets, contact me at ianstewartgmat at gmail.com

ianstewartgmat.com
Join the discussion

by winniethepooh » Wed Jun 15, 2011 4:32 pm
Thanks Ian!
Join the discussion

by edvhou812 » Wed Jun 15, 2011 9:05 pm
Probability Event A or B will happen is A+B - A*B.

Dice is 3: 1/6
Coin is Heads: 1/2

1/6+1/2 - 1/6*1/2

8/12-1/12 = 7/12
Join the discussion

by Ashley@VeritasPrep » Thu Jun 16, 2011 11:24 am
Diagram as promised! Note that this applies to the P(A) + P(B) - P(A-and-B) method, not to the alternative (also good!) method Ian proposes (1 - P(neither).
Image
You do not have the required permissions to view the files attached to this post.
Ashley Newman-Owens
GMAT Instructor
Veritas Prep

Post helpful? Mosey your cursor on over to that Thank button and click, please! I will bake you an imaginary cake.
Join the discussion